6x-2x+11=-5
4/3x-10/3x-11=1

Answer by checkley71(8403)

6X-2X+11=-5 OR 4X=-16 OR X=-16/4 OR X=-4
4/3X-10/3X-11=1 OR -6/3X=12 OR -2/X=12 OR 12X=-2 OR X=-1/6